# Equivariant CM minimization for extremal manifolds

### v1

@@ 3.4. Stability results
This is the relative version of [Hat24, Theorem 2.18].
@@
We come to the main result of this subsection: for this, assume that the torus $T$ in $\mathrm{Aut}(Y,H)$ is maximal.
@@
If $(X,L)$ is extremal, then for any $T$-invariant filtration $\mathcal{F}$ of $(X,L)$ with $||\mathcal{F}||_{T}>0$, we have
@@
for $T$-invariant filtrations $\mathcal{F}$ with $\langle\mathcal{F},N_{\mathbb{Z}}(T)\rangle=0$ and $||\mathcal{F}||>0$. This may be done by replacing $\mathcal{F}$ with the twist $\mathcal{F}_{-\xi(\mathcal{F})}$ for $\xi(\mathcal{F)}$ in $N_{\mathbb{R}}(T)$ as defined in Lemma 2.20. There is a slight subtlety that $\mathcal{F}_{-\xi(\mathcal{F})}$ is an $\mathbb{R}$-filtration, in general; we ignore this and hence assume that $\xi(\mathcal{F})$ is rational, as the argument applies in the same way in general, and $\xi(\mathcal{F})$ is rational in our applications.
Now, assume by contradiction that $\mathrm{Chow}_{\infty}(\mathcal{F})=0$. As $||\mathcal{F}||>0$, the essential infimum $M_{\mathcal{F}}$ of the concave transform $G_{\mathcal{F}}$ is strictly smaller than its average $\overline{G}_{\mathcal{F}}$ over $P$. Note that the values $M_{\mathcal{F}}$ and $\overline{G}_{\mathcal{F}}$ are both independent of choice of Okounkov body. Let
(3.6) $\Lambda_{\mathcal{F}}=\frac{n-2}{n-1}\overline{G}_{\mathcal{F}}+\frac{1}{n-1}M_{\mathcal{F}},$
which will serve as the value in Corollary 3.14 to obtain an upper bound for the concave transform near the origin. Indeed, since $M_{\mathcal{F}}<\overline{G}_{\mathcal{F}}$, we have $\Lambda_{\mathcal{F}}<e_{\mathrm{max}}(\mathcal{F})$ so by Corollary 3.14, there exists a $T$-invariant point $p$ in $X$ such that, after choosing an Okounkov body $P$ centered around $p$, there is an integer $m_{0}\geq 1$, such that
$G_{\mathcal{F}}\leq\Lambda_{\mathcal{F}}\quad\text{on}\quad\Delta_{\frac{1}{m_{0}}}\subset P.$
By assumption, for any $\delta>0$ and $s_{1}\geq 1$, there exists an integer $s\geq s_{1}$, such that $\mathrm{Chow}_{s}(\mathcal{F}_{(s)})<\delta$. We will proceed to set up appropriate choices of $\delta$ and $s_{1}$: let $\delta$ be a real number satisfying
$0<\delta<\frac{\overline{G}_{\mathcal{F}}-M_{\mathcal{F}}}{(n+3)(n-1)}.$
By [Szé15, Lemma 7], there exists an integer $s_{0}\geq 1$ such that for all $s\geq 0$, we have $\overline{G}_{\mathcal{F}}-\delta\leq\overline{G}_{\mathcal{F}_{(s)}}\leq\overline{G}_{\mathcal{F}}$. This gives rise the a priori estimates
$\int_{\partial_{0}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}}$ $\leq\frac{nm^{1-n}}{(n-1)!}\Lambda_{\mathcal{F}},$
$\int_{\partial_{1}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}}$ $\geq\frac{\sqrt{n}m^{1-n}}{(n-1)!}(M_{\mathcal{F}}-\delta),$
using that the boundary volumes of $\partial_{0}\Delta_{l}$ and $\partial_{0}\Delta_{l}$ are $nl^{n-1}$ and $\sqrt{n}l^{n-1}$, respectively. Moreover, by [Szé15, Lemma 7], the integral formula of the inner product, the boundedness of the weights and orthogonality of $\mathcal{F}$ with respect to $N_{\mathbb{Z}}(T)$, there exists an integer $s_{1}\geq s_{0}$, such that for all $s\geq 0$ and $m\geq m_{0}$, we have
@@
$\bigg{|}\sum_{i=1}^{d}\frac{\langle\mathcal{F}_{(s)},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\Big{(}\nu_{m}(\beta_{i})-\frac{b_{i,0}}{a_{0}}\Big{)}\bigg{|}$ $\leq\frac{a_{0}\delta}{n-1}.$
With these choices of $\delta$ and $s_{1}$, there exists an integer $s\geq s_{0}$ such that $\mathrm{Chow}_{s}(\mathcal{F}_{(s)})<\delta$. Let $\mathcal{F}_{(s)}^{\prime}$ be the $\overline{G}_{\mathcal{F}_{(s)}}$-shifted filtration of $\mathcal{F}_{(s)}$, given by
$\mathcal{F}_{(s)}^{\prime\lambda}H^{0}(Y,ksH)=\mathcal{F}_{(s)}^{\lambda+\overline{G}_{\mathcal{F}_{(s)}}k}H^{0}(Y,ksH).$
This filtration has the same Chow weights, while the leading coefficient of its weight function vanishes. In particular, the asymptotic expansion of Proposition 3.8 for this filtration takes the form
$\mathrm{Chow}_{r,T}(\widehat{\mathcal{F}_{(s)}^{\prime}})\leq 2\delta m^{-n}+\nu_{m,T}(p,\mathcal{F}_{(s)}^{\prime})\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0}).$
We now prove that the relative $\nu$-weight in this case is negative: write $\lambda=\frac{\overline{G}_{\mathcal{F}}-M_{\mathcal{F}}}{n-1}>0$, which satisfies both $\Lambda_{\mathcal{F}}=\overline{G}_{\mathcal{F}}-\lambda$ and $M_{\mathcal{F}}=\overline{G}_{\mathcal{F}}-(n-1)\lambda$. Using that $G_{\mathcal{F}_{(s)}^{\prime}}=G_{\mathcal{F}_{(s)}}-\overline{G}_{\mathcal{F}_{(s)}}$, the integrals appearing in the weight satisfy the estimates
$\int_{\partial_{0}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}^{\prime}}$ $\leq\frac{nm^{1-n}}{(n-1)!}(\delta-\lambda),$
$\int_{\partial_{1}\Delta_{\frac{1}{m}}}G_{\mathcal{F}_{(s)}^{\prime}}$ $\geq-\frac{\sqrt{n}m^{1-n}}{(n-1)!}(2\delta+(n-1)\lambda).$
Putting this into the $\nu$-weight, we obtain
$\nu_{m,T}(p,\mathcal{F}_{(s)}^{\prime})\leq\frac{a_{0}}{n-1}((n+3)\delta-\lambda),$
noting that the right-hand side is negative as $\delta<\frac{\lambda}{n+3}$ by construction.
Altogether, writing $\eta=\frac{\lambda-(n+3)\delta}{2(n-2)!n-1}>0$, we conclude
(3.7) $\mathrm{Chow}_{r,\widehat{T}}(\widehat{\mathcal{F}_{(s)}^{\prime}})\leq(2\delta-m\eta)m^{-n}+O(m^{1-n}r^{-1}).$
Now, choose $m\geq m_{1}$ to be large enough such $02\delta-m\eta<$ 0 and the blowup $(\widehat{Y},\widehat{sH}_{m})=(\mathrm{Bl}_{p}Y,msH-E)$ is extremal [APS11, Theorem 2.1]. But this is impossible, as [Sey17, Corollary 1.2] implies that the blowup is asymptotically relatively Chow semistable, but Equation (3.7) is negative for infinitely many $r$, contradicting Lemma 3.7. ∎
@@ 4.1. Models and filtrations
We start by recalling the construction of a good filtration from two models, due to Hattori [Hat24]. Let $R$ be a discrete valuation ring with fraction field $K$ and residue field $\mathbb{C}$ and let $(X,L)$ be a polarized $K$-variety. Let $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ be models of $(X,L)$ and assume that $L$, $\mathcal{L}$ and $\mathcal{L}^{\prime}$ are genuine line bundles.
[…]
@@
The main point of Hattori’s filtration is the following property:
@@
for $k\gg 0$. Moreover $||\mathcal{F}||>0$ if and only if the models are not isomorphic.
@@
For the norm of the filtration, it suffices to show that if the difference $\mu^{*}\mathcal{L}-\mathcal{M}$ supported in $\mathcal{Y}_{0}$ is not $\mathbb{Q}$-linearly trivial, then the concave transform $G_{\mathcal{F}}$ is not constant. Similar to Section 2.1.1, we may assume that $E$ is non-zero, $\mu$-exceptional and $-E$ is $\mu$-ample by taking an ample model.
@@
which is strictly positive for all $10<s<$ 1 by ampleness. Since the function satisfies $f(0)=0$ and $f(1)=\mu^{*}\mathcal{L}^{n+1}-\mathcal{M}^{n+1}$, we thus obtain
@@
=$=\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(k\mathcal{L})\text{ and }t^{-\lambda}s\in i_{\beta}(H^{0}(kr\mathcal{M}))\},$
@@
We have $||\mathcal{F}||_{T_{0}}>0$ if and only if $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all $\beta$ one-parameter subgroups $\beta$ of $T$.
@@
$||\mathcal{F}||_{T_{0}}=\min_{\xi\in N_{\mathbb{Q}}(T)}||\mathcal{F}_{\xi}||,$
where the twist $\mathcal{F}_{\xi}$ may be viewed as a $\mathbb{Z}$-filtration of $(\mathcal{X}_{0},r\mathcal{L}_{0})$ for a sufficiently divisible positive integer $r$. Applying Theorem 4.2, we obtain $||\mathcal{F}||_{T_{0}}>0$ if and only if $(\mathcal{X}_{\xi},\mathcal{L}_{\xi})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all rational twists $\xi$ in $N_{\mathbb{Q}}(T)$. But this is equivalent to $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all $\beta$ in $N_{\mathbb{Z}}(T)$ as we may simply scale a given automorphism. ∎
@@ 4.3. Proof of the main results
for $\xi=-\sum_{i=1}^{d}\langle\mathcal{F},\beta_{i}\rangle/\langle\beta_{i},\beta_{i}\rangle\beta_{i}$, where $(\beta_{i})$ is any of $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$. As $(\mathcal{X}_{0},\mathcal{L}_{0})$ is extremal, we have $\mathrm{DF}_{T_{0}}(\mathcal{F})\geq 0$ with strict inequality if and only if $||\mathcal{F}||_{T_{0}}>0$ by Theorem 3.2. The result follows by Proposition 4.6. ∎
@@
The proof reveals that we may explicitly give the rational one-parameter subgroup $\xi$ as $-\xi(\mathcal{F})=-\sum_{i=1}^{d}\langle\mathcal{F},\beta_{i}\rangle/\langle\beta_{i},\beta_{i}\rangle\beta_{i}$ from Lemma 2.20. Geometrically, it may be viewed as the orthogonal projection of the filtration $\mathcal{F}$ to $N_{\mathbb{Q}}(T)$ whose coefficients may be expressed geometrically in terms of Hilbert- and weight polynomials of the models, following Theorem (4.2) and Lemma 4.3.

### v2

@@ 3.4. Stability results
This is a relative version of [Hat24, Theorem 2.18].
@@
We come to the main result of this subsection. For this, assume that the torus $T$ in $\mathrm{Aut}(Y,H)$ is maximal.
@@
If $(X,L)$ is extremal, then for any $T$-invariant filtration $\mathcal{F}$ of $(X,L)$ with $\mathcal{F}||_{T}>0$, we have
@@
for $T$-invariant filtrations $\mathcal{F}$ with $\langle\mathcal{F},N_{\mathbb{Z}}(T)\rangle=0$ and $\mathcal{F}||>0$. This may be done by replacing $\mathcal{F}$ with the twist $\mathcal{F}_{-\xi(\mathcal{F})}$ for $\xi(\mathcal{F)}$ in $N_{\mathbb{R}}(T)$ as defined in Lemma 2.21. There is a slight subtlety that $\mathcal{F}_{-\xi(\mathcal{F})}$ is an $\mathbb{R}$-filtration, in general; we ignore this and hence assume that $\xi(\mathcal{F})$ is rational, as the argument applies in the same way in general, and $\xi(\mathcal{F})$ is rational in our applications.
Now, assume by contradiction that $\mathrm{Chow}_{\infty}(\mathcal{F})=0$. As $\mathcal{F}||>0$, the essential infimum $M_{\mathcal{F}}$ of the concave transform $G_{\mathcal{F}}$ is strictly smaller than its average $\overline{G}_{\mathcal{F}}$ over $P$. Note that the values $M_{\mathcal{F}}$ and $\overline{G}_{\mathcal{F}}$ are both independent of choice of Okounkov body. Let
(3.5) $\Lambda_{\mathcal{F}}=\frac{A_{n}-2}{A_{n}}\overline{G}_{\mathcal{F}}+\frac{1}{A_{n}}M_{\mathcal{F}},$
for $A_{n}=\frac{5n-1}{2n}$. This number will serve as the value in Corollary 3.14 to obtain an upper bound for the concave transform near the origin. Indeed, since $M_{\mathcal{F}}<\overline{G}_{\mathcal{F}}$, we have $\Lambda_{\mathcal{F}}<e_{\mathrm{max}}(\mathcal{F})$ so by Corollary 3.14, there exists a $T$-invariant point $p$ in $X$ such that, after choosing an Okounkov body $P$ centered around $p$, there is an integer $m_{0}\geq 1$, such that
$G_{\mathcal{F}}|_{\Delta_{\frac{1}{m_{0}}}}\leq\Lambda_{\mathcal{F}}.$
By assumption, for any $\delta>0$ and $s_{1}\geq 1$, there exist infinitely many integers $s\geq s_{1}$ such that $\mathrm{Chow}_{1}(\mathcal{F}_{(s)})<\delta$. We will proceed to set up appropriate choices of $\delta$ and $s_{1}$: let $\delta$ be a real number satisfying
$0<\delta<\frac{\overline{G}_{\mathcal{F}}-M_{\mathcal{F}}}{A_{n}B_{n}},$
where $B_{n}=\frac{7n-1}{2n}$. By [Szé15, Lemma 7], there exists an integer $s_{0}\geq 1$ such that for all $s\geq s_{0}$, we have $\overline{G}_{\mathcal{F}}-\delta\leq\overline{G}_{\mathcal{F}_{(s)}}\leq\overline{G}_{\mathcal{F}}$. This gives rise to an a priori estimate
$\int_{\Delta_{\frac{1}{m}}\setminus\Delta_{\frac{1}{m}-\frac{n}{r}}}G_{\mathcal{F}_{(s)}}\geq\frac{nm^{1-n}}{r(n-1)!}(M_{\mathcal{F}}-\delta),$
for any $r\geq 1$. Moreover, by [Szé15, Lemma 7], the integral formula of the inner product, the boundedness of the weights and orthogonality of $\mathcal{F}$ with respect to $N_{\mathbb{Z}}(T)$, there exists an integer $s_{1}\geq s_{0}$, such that for all $s\geq 0$ and $m\geq m_{0}$, we have
@@
$\bigg|\sum_{i=1}^{d}\frac{\langle\mathcal{F}_{(s)},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\Big(\nu(\beta_{i})-\frac{b_{i,0}}{a_{0}}\Big)\bigg|$ $\leq\frac{a_{0}\delta}{n-1},$
where $a_{0}$ is from Equation 2.5. With these choices of $\delta$ and $s_{1}$, there exist infinitely many integers $s\geq s_{0}$ such that $\mathrm{Chow}_{1}(\mathcal{F}_{(s)})<\delta$. Let $\mathcal{F}^{\prime}$ be the $\overline{G}_{\mathcal{F}}$-shifted filtration of $\mathcal{F}$, given as
$\mathcal{F}^{\prime\,\lambda}H^{0}(Y,kH)=\mathcal{F}^{\lambda+\overline{G}_{\mathcal{F}}k}H^{0}(Y,kH).$
This filtration has the same Chow weights, while the leading coefficient of its weight function vanishes. By taking $s$ to be large enough and changing $m$, we may assume that $\frac{s}{m}=:r$ is an integer.
The relative asymptotic expansion of Theorem 3.8 for this filtration then takes the form
$\mathrm{Chow}_{r,T}(\widehat{\mathcal{F}^{\prime}})\leq 2\delta+\nu_{T}(p,\mathcal{F}_{(s)}^{\prime})\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0}).$
We now prove that the $\nu_{T}$-weight is negative: write $\gamma_{n}=\frac{\overline{G}_{\mathcal{F}}-M_{\mathcal{F}}}{a_{n}}>0$, which satisfies both $\Lambda_{\mathcal{F}}=\overline{G}_{\mathcal{F}}-\gamma_{n}$ and $M_{\mathcal{F}}=\overline{G}_{\mathcal{F}}-A_{n}\gamma_{n}$. Using the relation $G_{\mathcal{F}_{(s)}^{\prime}}=G_{\mathcal{F}_{(s)}}-\overline{G}_{\mathcal{F}_{(s)}}$, the summands appearing in the weight satisfy the estimates
$\frac{(3n-1)}{2(n-1)!}\Lambda_{\mathcal{F}^{\prime}_{(s)}}$ $\leq\frac{n}{(n-1)!}\frac{3n-1}{2n}(\delta-\gamma_{n}),$
$\frac{r}{m^{n-1}}\int_{\Delta_{\frac{1}{m}}\setminus\Delta_{\frac{1}{m}-\frac{n}{r}}}G_{\mathcal{F}^{\prime}_{(s)}}$ $\geq-\frac{n}{(n-1)!}(\delta+A_{n}\gamma_{n}).$
Using that $G_{\mathcal{F}^{\prime}_{(s)}}\leq(\delta-\gamma_{n})$ on $\Delta_{\frac{1}{m_{0}}}$. Putting this into the weight, we obtain
$\nu_{T}(p,\mathcal{F}_{(s)}^{\prime})\leq\frac{a_{0}}{n-1}(b_{n}\delta m^{-1}-\gamma_{n}),$
noting that the right-hand side is negative as $\delta<\frac{\gamma_{n}}{b_{n}}$ by construction.
Altogether, we conclude that
(3.6) $\mathrm{Chow}_{r,\widehat{T}}(\widehat{\mathcal{F}^{\prime}})\leq\Big(2\delta+\frac{a_{0}}{n-1}(b_{n}\delta-m\gamma_{n})m^{-n}\Big)+O(m^{1-n}r^{-1}).$
Now, choose $m\geq m_{1}$ to be large enough such the leading term above is negative and the blowup $(\widehat{Y},\widehat{sH}_{m})=(\mathrm{Bl}_{p}Y,msH-E)$ is extremal [APS11, Theorem 2.1]. But this is impossible, as [Sey17, Corollary 1.2] implies that the blowup is asymptotically relatively Chow semistable, but Equation (3.6) is negative for infinitely many $r$, contradicting Lemma 3.7. ∎
@@ 4.1. Models and filtrations
We start by giving the construction of a good filtration from two models, due to Hattori [Hat24]. Let $R$ be a discrete valuation ring with fraction field $K$ and residue field $\mathbb{C}$ and let $(X,L)$ be a polarized $K$-variety. Let $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ be models of $(X,L)$ and assume that $L$, $\mathcal{L}$ and $\mathcal{L}^{\prime}$ are ($\mathbb{Z}$-)line bundles.
We repeat the construction in Equation (2.1) by letting $\mathcal{X}^{\prime}\xleftarrow{\mu^{\prime}}\mathcal{Y}\xrightarrow{\mu}\mathcal{X}$ be a resolution of indeterminacies of the canonical birational map $\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}$ over $\operatorname{Spec}R$, and write $\mu^{*}\mathcal{L}=\mu^{\prime*}\mathcal{L}^{\prime}+E-a\mathcal{Y}_{0}$ for an effective divisor $E$ in $\mathcal{Y}_{0}$ and an integer $a\geq 1$. We obtain a natural embedding
@@
The main property of Hattori’s filtration is the following:
@@
for $k\gg 0$. Moreover $\mathcal{F}||>0$ if and only if the models are not isomorphic.
@@
For the norm of the filtration, it suffices to show that if the difference $\mu^{*}\mathcal{L}$-$\mathcal{M}$ supported in $\mathcal{Y}_{0}$ is not $\mathbb{Q}$-linearly trivial, then the concave transform $G_{\mathcal{F}}$ is not constant. Similar to Section 2.1.1, we may assume that $E$ is non-zero, $\mu$-exceptional and E-E is $\mu$-ample by taking an ample model.
@@
which is strictly positive for all $0<s<1$ by ampleness. Since the function satisfies $f(0)=0$ and $f(1)=\mu^{*}\mathcal{L}^{n+1}-\mathcal{M}^{n+1}$, we thus obtain
@@
$=\{s|_{\mathcal{X}_{0}}\;|\;s\in H^{0}(k\mathcal{L})\text{ and }t^{-\lambda}s\in i_{\beta}(H^{0}(kr\mathcal{M}))\},$
@@
We have $\mathcal{F}||_{T_{0}}>0$ if and only if $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all $\beta$ one-parameter subgroups $\beta$ of $T$.
@@
[…]
where the twist $\mathcal{F}_{\xi}$ may be viewed as a $\mathbb{Z}$-filtration of $(\mathcal{X}_{0},r\mathcal{L}_{0})$ for a sufficiently divisible positive integer $r$. Applying Theorem 4.2, we obtain $\mathcal{F}||_{T_{0}}>0$ if and only if $(\mathcal{X}_{\xi},\mathcal{L}_{\xi})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all rational twists $\xi$ in $N_{\mathbb{Q}}(T)$. But this is equivalent to $(\mathcal{X}_{\beta},\mathcal{L}_{\beta})\not\cong(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ for all $\beta$ in $N_{\mathbb{Z}}(T)$ as we may simply scale a given automorphism. ∎
@@ 4.3. Proof of the main results
for $\xi=-\sum_{i=1}^{d}\langle\mathcal{F},\beta_{i}\rangle/\langle\beta_{i},\beta_{i}\rangle\beta_{i}$, where $(\beta_{i})$ is any of $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$. As $(\mathcal{X}_{0},\mathcal{L}_{0})$ is extremal, we have $\mathrm{DF}_{T_{0}}(\mathcal{F})\geq 0$ with strict inequality if and only if $\mathcal{F}||_{T_{0}}>0$ by Theorem 3.2. The result follows by Proposition 4.6. ∎
@@
The proof reveals that we may explicitly give the rational one-parameter subgroup $\xi$ as $\xi(\mathcal{F})=-\sum_{i=1}^{d}\langle\mathcal{F},\beta_{i}\rangle/\langle\beta_{i},\beta_{i}\rangle\beta_{i}$ from Lemma 2.21. Geometrically, it may be viewed as the orthogonal projection of the filtration $\mathcal{F}$ to $N_{\mathbb{Q}}(T)$ whose coefficients may be expressed geometrically in terms of Hilbert- and weight polynomials of the models, following Theorem (4.2) and Lemma 4.3.